# Towards Next Era of Multi-objective Optimization: Large Language Models as Architects of Evolutionary Operators

### v1

@@ Abstract
Multi-objective optimization problems (MOPs) are prevalent in various real-world applications, necessitating sophisticated solutions that balance conflicting objectives. Traditional evolutionary algorithms (EAs), while effective, often rely on domain-specific expert knowledge and iterative tuning, which can impede innovation when encountering novel MOPs. Very recently, the emergence of Large Language Models (LLMs) has revolutionized software engineering by enabling the autonomous development and refinement of programs. Capitalizing on this advancement, we propose a new LLM-based framework for evolving EA operators, designed to address a wide array of MOPs. This framework facilitates the production of EA operators without the extensive demands for expert intervention, thereby streamlining the design process. To validate the efficacy of our approach, we have conducted extensive empirical studies across various categories of MOPs. The results demonstrate the robustness and superior performance of our LLM-evolved operators.
@@ I Introduction
[lines=2]Multi-objective optimization is a crucial field of research that tackles the challenge of simultaneously optimizing two or more conflicting objectives in real-world scenarios[1]. The primary goal of multi-objective optimization is to find a set of solutions, often known as the Pareto front [2], where no other solutions outperform all objectives. To solve the multi-objective optimization problems (MOPs), a lot of well-known evolutionary algorithms (EAs) have been developed in the literature, including non-dominated-sorting based methods [3, 4, 5] and decomposition based techniques [6, 7, 8]. With the growing demand of multi-objective optimization in real-world applications [9], extensive efforts have been conducted to improve the search performance of these methods, with the goal of tackling the new multi-objective optimization challenges encountered. However, such hand-crafting pipeline for algorithm design necessitates a substantial amount of expert knowledge and numerous trials, which constrains the innovation of EAs.
Nowadays, owing to the impressive language processing capabilities, Large Language Models (LLMs) have made significant strides in software engineering tasks, as evidenced by recent works [10, 11]. Particularly, Chen et al. enabled LLMs to refine their programs autonomously [12]; Shypula et al. proposed to enhance the efficiency of programs via few-shot prompting and chain-of-thought [13]. Moreover, peoples have been used to generate their own programs to fit the daily programming demands by leveraging powerful LLMs such as ChatGPT [14]. However, the creation of effective programs with sophisticated functionalities requires more than just a handful of simple human-machine interaction. Building on the concept of iterative evolution, Romera-Paredes et al. utilized LLMs to make new mathematical discoveries via search novel programs iteratively, indicating the potential software engineering capability by the integration of LLMs and evolutionary computation [15]. Liventsev et al. have showcased the performance of fully autonomous programming using LLMs, suggesting a future where programs evolve without human guidance [16]. Liu et al. proposed a LLM-based algorithm evolution framework to design novel programs for solving traveling salesman problems (TSPs) [17]. Subsequently, they extended their work to develop superior guided local search for combinatorial optimization [18]. Furthermore, Pluhacek et al. leveraged LLMs in meta-heuristic optimization to generate hybrid swarm intelligence, potentially outperforming traditional methods [19]. Ye et al. introduced LLMs as hyper-heuristics within a reflective evolution framework, allowing for dynamic adaptation [20]. As these works demonstrate, LLMs are revolutionizing the field of software engineering through iterative program evolution, marking a new era in algorithmic design for optimization.
While the existing methods of program evolution have achieved notable successes in various domains, it should be noted that, they predominantly focus on the exploitation of pivotal task-specific heuristics using succinct prompts [21]—for example, guided local search techniques [18]—to enhance the search efficacy on particular problems [18, 19, 20]. However, these novel heuristics often face challenges when applied to diverse and complex issues. Additionally, their application in designing programs with intricate functionalities frequently leads to catastrophic outcomes, such as program crashes or non-terminating processes, primarily due to inadequate error-handling mechanisms.
@@
With these considerations in mind, we propose a new operator evolution framework that utilizes LLMs to discover and refine EA operators tailored for diverse MOPs. Recognizing the susceptibility to errors and the prevalence of execution anomalies in sophisticated programs produced by LLMs, our framework incorporates a robust testing module, which is seamlessly integrated into the evolutionary progress of the EA operators. This module treats compilation and runtime errors as indicators, and then guides the refinement of code through dialogues with LLMs using the collected errors. To safeguard against the risk of infinite execution loops—a known hazard in LLM-generated code—we launch this module as an independent process, governed by a strict execution time limit. Moreover, we developed a dynamic selection module to cultivate a variety of EA operators. This diversity is key to circumventing premature convergence to local optima, thereby enhancing the exploration capabilities of the prompting-based evolutionary process.
Particularly, our proposed framework initiates with a diverse population of EA operators generated through LLMs. Throughout the evolutionary process, our proposed textual crossover and mutation techniques guide the LLM to craft advanced EA operators for various MOPs. These EA operators generated via LLMs may encompass various strategies such as “selection”, “crossover”, “mutation” to effectively address MOPs. The contributions of this work can be outlined as follows:
@@
  • To handle the challenges posed by multi-objective optimization, a new EA operator evolution framework is devised, which facilitates the autonomous ideation and crafting of cutting-edge EA operators, thereby amplifying the search power for MOPs.
@@
  • In light of the intrinsic vulnerabilities and the frequently occurred errors in programs conceived by LLMs, a robust testing module is established, which can seamlessly refine the generated code by leveraging the errors as a dialogue-based feedback with LLMs.
@@
  • To enhance the performance of EA operator evolution, we introduce a dynamic selection strategy tailored to LLM-based textual crossover module.
@@
  • To validate the performance of our proposed framework, we have embarked on extensive empirical studies, employing both continuous and combinatorial MOPs, against human-engineered multi-objective methodologies. The obtained results demonstrate the performance of EA operators generated via the proposed framework.
@@ II Background
This section begins with a literature review of the LLM-assisted program evolution methods. Subsequently, we introduce three specific MOP categories: Continuous Multi-objective Optimization Problems (CMOPs), Multi-objective Knapsack Problems (MOKPs), and Multi-objective Traveling Salesman Problems (MOTSPs), serving as the focal points of investigation in our study. Notably, the rationale behind the selection of CMOP, MOKP and MOTSP is to investigate the capability of LLM in devising innovative EA operators across varied gene encoding modes.
II-A Literature Review
The advent of LLMs has ushered in a transformative era in program evolution, enabling the generation of code that caters to diverse programming demands [10, 11]. Early study of Lehman et al. explored the potential of LLM trained to generate code in the context of genetic programming, which reveals that LLMs can significantly enhance the effectiveness of the programs. Subsequently, Romera-Paredes et al. [15] and Liventsev et al. [16] demonstrated the potential of LLMs in iterative evolution and fully autonomous programming, respectively, suggesting a future where human intervention in code generation may become minimal. Furthermore, Liu et al.’s LLM-based algorithm evolution framework has shown promise in solving combinatorial optimization problems such as the traveling salesman problem [17], while Pluhacek et al. have utilized LLMs to enhance meta-heuristic optimization through hybrid swarm intelligence [19]. The reflective evolution framework introduced by Ye et al. leverages LLMs as hyper-heuristics, allowing for dynamic adaptation to evolving problem landscapes [20].
These advancements underscore the transformative role of LLMs in the algorithmic design for optimization tasks. Despite the promising potential, the exploration of LLM-assisted program evolution in optimization remains embryonic. To date, no initiatives have been undertaken to develop innovative multi-objective optimization methods via LLMs. This gap highlights the imperative need for our proposed methodology, which aims to overcome the intrinsic constraints of LLMs in crafting evolutionary operators for superior multi-objective optimization outcomes.
II-B Multi-objective Optimization Problems
II-B1 Continuous Multi-objective Optimization Problems
CMOPs represent a critical class of optimization challenges [22]. These problems involve optimizing multiple conflicting objectives simultaneously, where the decision variables are continuous real numbers. Considering a minimization CMOP, the mathematical definition can be given by:
@@ II-B1 Continuous Multi-objective Optimization Problems
where $m$ represents the number of objectives, $\mathbf{x}$ denotes the decision variables and $\mathcal{X}$ specifies the feasible search space. Generally, the goal of CMOP is to find a set of Pareto-optimal solutions, $\mathcal{P}=\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots\},\mathbf{x}\in\mathcal{X}$, which cannot be improved in any objective without degrading another. The research value of CMOPs lies in their applicability to real-world scenarios, such as engineering design [23], finance [24], and environmental management [25].
II-B2 Multi-objective Knapsack Problems
@@
MOKPs find applications in many real-world optimization problems, MOKPs find applications in many real-world optimization problems, such as food order optimization [27], engineering and operational research [28], cargo loading and project selection [29]. The versatility of MOKPs in addressing these diverse challenges underscores their significance in strategic planning and optimization across various industries.
II-B3 Multi-objective Traveling Salesman Problems
@@
III Operator Evolution for MOPs
In this section, the proposed paradigm for evolving EA operators via LLM is introduced. The workflow of our proposed paradigm has been depicted in Fig. 1. In contrast to existing program evolution pipelines [34, 17, 15], our proposed method integrates an LLM-assisted debugging process, encompassing Pilot Run & code repair, into the core operator generation stages: “Operator Initialization”, “Operator Crossover”, and “Operator Mutation”. Additionally, the “Dynamic Operator Selection” is tailored to complement the LLM-based crossover mechanism, facilitating the handling of textual data. As illustrated in Fig. 1 and outlined in Alg. 1, the paradigm begins with the phase of “Operator Initialization”, where an initial population of EA operators is generated from a meticulously crafted ‘Initial Prompts’ (line 1 of Alg. 1). These operators undergo “Parallel Operator Evaluation”, receiving scores based on their efficacy across MOPs. These scores inform the “Dynamic Operator Selection” (lines 4-5), promoting diversity in “Crossover Prompts” by altering the number of selected parent operators, thus fostering the creation of innovative operators during “Operator Crossover” (line 6) and “Operator Mutation” (line 7), based on dialogues with LLM. The newly generated operator obtained through mutation (i.e., $\hat{RC}$) is evaluated in parallel to obtain its score, which is then integrated into the population, followed by an elitist strategy-based update of the population (lines 8-11). The evolution of operators persists, driven by enhanced optimization performance, until the termination criteria are met.
@@ III Operator Evolution for MOPs
5             Generate a random integer $N_{s}$, $1<N_{s}<N_{ev}$
@@ 10       $\mathcal{P}_{ev}=\mathcal{P}_{ev}\cup\{RC\}$
The prompts for initialization are presented in the subsequent box. As depicted, the prompts can be briefly divided into three parts, i.e., ‘System’, ‘Description of Task’ and ‘Requirements’, respectively. The ‘System’ gives the initial prompts for LLM with a brief description for the task. The ‘Description of Task’ outlines the details of the task for the LLM, while the ‘Requirements’ provides the specific requirements on the task. Within these prompts, the placeholder #PROBLEM# is a user-defined descriptor for MOPs, such as the ‘multi-objective knapsack problem’. The #PROBLEM_DESC# presents an informative introduction to the MOP with optimization objectives, which aids crafting bespoke search operator for specific MOPs. Additionally, #FORMAT# delineates the operator’s specifics, encompassing the function name, input parameters, output and their annotations. Details of these placeholders have been given in the Supplementary Materials. Employing these comprehensive prompts enables the LLM to generate EA operators that may achieve higher score across MOPs, as delineated in Alg. 2. To further validate the operator produced by the LLM engine, it will undergo Pilot Run & Repair as discussed in Section III-E, specifically lines 3-7 of Alg. 2. The generation process continues until an operator successfully passes the validation tests or a rectified operator is produced through the repair process outlined in Alg. 5.
@@ III-B Dynamic Operator Selection via Scores
Considering the aforementioned issue, this study introduces a dynamic selection mechanism to facilitate the generation of novel operators during operator crossover. Specifically, as illustrated in lines 4-5 of Algorithm 1, the selection probabilities for the individual operators within $\mathcal{P}_{ev}$ are determined based on their performance scores on MOPs. A higher score signifies enhanced search capability and, consequently, an increased likelihood of selection. The selection probability for each operator is computed as follows:
@@
where $N_{ev}$ denotes the number of operators within the population. Additionally, a random integer $N_{s}$, where $21<N_{s}<N_{ev}/$ 2, is generated to designate the count of parent operators chosen for the operator crossover.
@@ III-C Operator Crossover
Utilizing the dynamically selected parent operators, this study performs diverse operator crossovers by engaging a LLM with these operators. The LLM, serving as an intelligent engine, is anticipated to synthesize innovative operators that surpass the parent operators in terms of scores. The prompts for the LLM are structured as follows:
@@
The placeholders #PROBLEM#, #FORMAT#, #$N_{s}$#, and #SELECTED_OPERATORS# denote the MOP name, the coding format of the operator, the quantity of parent operators selected, and the XML code snippets of the parent operators with their scores, respectively, which have been detailed in Supplementary Materials. The dynamic nature of these prompts is likely to facilitate the generation of varied operators by the LLM, as outlined in line 3 of Alg. 3. However, as outlined in line 4 of Alg. 1, the resultant operator $TC$ requires additional validation or refinement to obtain $RC$, akin to the proposed operator initialization.

### v2

# Autonomous Multi-Objective Optimization Using Large Language Model

@@ Abstract
Multi-objective optimization problems (MOPs) are ubiquitous in real-world applications, presenting a complex challenge of balancing multiple conflicting objectives. Traditional evolutionary algorithms (EAs), though effective, often rely on domain-specific expertise and iterative fine-tuning, hindering adaptability to unseen MOPs. In recent years, the advent of Large Language Models (LLMs) has revolutionized software engineering by enabling the autonomous generation and refinement of programs. Leveraging this breakthrough, we propose a new LLM-based framework that autonomously designs EA operators for solving MOPs. The proposed framework includes a robust testing module to refine the generated EA operator through error-driven dialogue with LLMs, a dynamic selection strategy along with informative prompting-based crossover and mutation to fit textual optimization pipeline. Our approach facilitates the design of EA operators without the extensive demands for expert intervention, thereby speeding up the innovation of EA operators. Empirical studies across various MOP categories validate the robustness and superior performance of our proposed framework.
@@ I Introduction
[lines=2]Multi-objective optimization is a crucial field of research that tackles the challenge of simultaneously optimizing two or more conflicting objectives in real-world scenarios[1]. The primary goal of multi-objective optimization is to find a set of solutions, often known as the Pareto Optimal Set [2], where no other solutions outperform all objectives. To solve the multi-objective optimization problems (MOPs), a lot of well-known evolutionary algorithms (EAs) have been developed in the literature, including non-dominated-sorting-based methods [3, 4, 5] and decomposition-based techniques [6, 7, 8]. With the growing demand of multi-objective optimization in real-world applications [9], extensive efforts have been conducted to improve the search performance of EAs [10, 11, 12], with the goal of tackling the new multi-objective optimization challenges encountered. However, such hand-crafting pipeline for algorithm design necessitates a substantial amount of expert knowledge and numerous trials, which could constrain the innovation of EA in solving complex optimization problems.
Nowadays, owing to the impressive language processing capabilities, Large Language Models (LLMs) have made significant strides in software engineering tasks. Modern society increasingly leverages these powerful LLMs to address the demands of daily programming tasks [13, 14, 15]. Particularly, Chen et al. enabled LLMs to refine their programs autonomously [16], while Shypula et al. proposed to enhance the efficiency of programs via few-shot prompting and chain-of-thought [17]. Furthermore, efforts to create powerful programs using LLMs have explored iterative evolution, leading to new mathematical discoveries [18], autonomous programming [19], enhanced vehicle routing [20, 21], hybrid swarm intelligence [22], and dynamic heuristic adaptation [23]. As these works demonstrate, LLMs are revolutionizing the field of software engineering through iterative program evolution, marking a new era in algorithmic design for optimization.
While the existing methods of program evolution have showcased notable successes across various domains, it is worth noting that, they often adopt succinct prompts and traditional evolutionary strategies. Such design effectively develops pivotal task-specific functions with locked power of LLMs [21, 22, 23]. However, when employed to construct intricate problem-solving systems for handling complex problems, such as MOPs, these methods often result in catastrophic outcomes—such as program crashes or endless loop—mainly due to inadequate error-handling mechanisms. Furthermore, they may encounter bottlenecks in program design arising from the absence of informative task descriptions and mismatched evolutionary strategies. Although these strategies perform well in numerical optimization tasks, they do not seamlessly align with the prompting-based textual optimization paradigm, which often lead to early stagnation of the search.
@@
With the above in mind, to further exploit the power of LLM, we propose a new framework of EA operator design that utilizes LLMs to discover and refine EA operators tailored for diverse MOPs. Recognizing the susceptibility to errors and the prevalence of execution anomalies in sophisticated programs produced by LLMs, our framework incorporates a robust testing module, which is seamlessly integrated into the evolutionary progress. This module treats compilation and runtime errors as indicators, and then guides the refinement of designed operator through dialogues with LLMs using the collected errors. To safeguard against the risk of infinite execution loops—a known hazard in LLM-generated code—we launch this module as an independent process, governed by a strict execution time limit. Furthermore, we introduce a dynamic selection module, along with informative crossover and mutation that align with the prompting-based textual optimization paradigm, aims to enlarge the exploration capabilities of LLMs. Particularly, our proposed framework initiates with a diverse population of EA operators generated through LLMs. Throughout the evolutionary process, our proposed textual crossover and mutation techniques guide the LLM to craft advanced EA operators for various MOPs. These EA operators generated via LLMs may encompass various search strategies to effectively address MOPs. The contributions of this work can be outlined as follows:
@@
  • To handle the challenges posed by autonomous multi-objective optimization, a new EA operator evolution framework is devised, which facilitates the autonomous ideation and crafting of cutting-edge EA operators, thereby amplifying the search efficacy for solving MOPs.
@@
  • In light of the intrinsic vulnerabilities and the frequently occurred errors in programs conceived by LLMs, a robust testing module is established, which can seamlessly refine the generated operator by leveraging the errors as a dialogue-based feedback with LLMs.
@@
  • To enhance the performance of EA operator evolution, we introduce a dynamic selection strategy, along with informative crossover and mutation specifically tailored to the LLM-based textual optimization paradigm.
@@
  • To validate the performance of our proposed framework, we have conducted extensive empirical studies, employing both continuous and combinatorial MOPs, against state-of-the-art human-engineered multi-objective optimization algorithms. The obtained results demonstrate the efficacy of EA operators generated via the proposed framework.
@@ II Background
This section begins with a literature review of the LLM-assisted program evolution methods. Subsequently, the mathematical definitions of continuous MOPs and combinatorial MOPs are given, which serve as the focal points of investigation in our study. Notably, the rationale behind the selection of these MOPs is to investigate the capability of LLM in devising innovative EA operators across varied gene encoding modes.
II-A Review of LLM-assisted Program Evolution
The emergence of LLMs have marked a transformative phase in optimization [14, 24, 25]. Within the context of textual optimization, such as program evolution, LLMs empower the generation of code that addresses a wide range of programming requirements [13]. Early study of Lehman et al. explored the potential of LLM trained to generate code in the context of genetic programming, which reveals that LLMs can significantly enhance the effectiveness of the programs [26]. Subsequently, Romera-Paredes et al. utilized LLMs to make new mathematical discoveries via search novel programs iteratively, indicating the potential software engineering capability by the integration of LLMs and evolutionary computation [18]. Liventsev et al. showcased the performance of fully autonomous programming using LLMs, suggesting a future where programs evolve without human guidance [19]. Liu et al. proposed a LLM-based algorithm evolution framework to design novel programs for solving traveling salesman problems (TSPs) [20]. Furthermore, they extended their work to develop superior guided local search for combinatorial optimization [21]. Pluhacek et al. leveraged LLMs in meta-heuristic optimization to generate hybrid swarm intelligence, potentially outperforming traditional methods [22]. Ye et al. introduced LLMs as hyper-heuristics within a reflective evolution framework, allowing for dynamic adaptation [23]. These advancements underscore the transformative role of LLMs in the algorithmic design for optimization tasks.
Despite the promising advancements, the exploration of LLM-assisted program evolution in optimization remains embryonic. To date, no effort has been made to develop innovative EA operators using LLMs for MOPs. Our proposed methodology aims to bridge this gap, overcoming the intrinsic constraints of LLMs and crafting EA operators that deliver superior search performance across different MOPs.
II-B Continuous Multi-objective Optimization Problems
Continuous Multi-objective Optimization Problems (CMOPs) constitute a pivotal category of optimization problems. These problems, characterized by continuous real numbers as decision variables, entail the optimization of multiple conflicting objectives. Considering a minimization CMOP, the mathematical definition can be given by:
@@ II-B Continuous Multi-objective Optimization Problems
where $k$ represents the number of objectives, $\mathbf{x}$ denotes the decision variables and $\mathcal{X}$ specifies the feasible search space. Generally, the goal of CMOP is to find a set of Pareto-optimal solutions, $\mathcal{P}=\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots\},\mathbf{x}\in\mathcal{X}$, which cannot be improved in any objective without degrading another. The research value of CMOPs lies in their applicability to real-world scenarios, such as engineering design [27], finance [28], and environmental management [29].
II-C Combinatorial Multi-objective Optimization Problems
In contrast to the continuous MOPs which involve real-valued decision variables, combinatorial MOPs focus on discrete decision variables, which present challenges due to their discontinuity and lack of smoothness. In what follows, we introduce two distinct combinatorial MOPs, each characterized by unique gene encoding.
II-C1 Multi-objective Knapsack Problems
@@
MOKPs find applications in many real-world optimization problems, such as food order optimization [31], engineering and operational research [32], cargo loading and project selection [33]. The versatility of MOKPs in addressing these diverse challenges underscores their significance in strategic planning and optimization across various industries.
II-C2 Multi-objective Traveling Salesman Problems
@@
III Proposed Method
@@
5             Generate a random integer $N_{s}$, $1<N_{s}\leq N_{max}$
@@
In this section, the proposed method for evolving EA operators via LLM is introduced. The workflow of our proposed paradigm has been depicted in Fig. 1. In contrast to existing program evolution pipelines [26, 20, 18], our proposed method integrates an LLM-assisted debugging process, encompassing Pilot Run & Repair, into the core operator generation stages: “Operator Initialization”, “Operator Crossover”, and “Operator Mutation”. Additionally, the “Dynamic Operator Selection” is tailored to complement the LLM-based crossover mechanism, facilitating the handling of textual data.
As illustrated in Fig. 1 and outlined in Alg. 1, the paradigm begins with the phase of “Operator Initialization”, where an initial population of EA operators is generated from a meticulously crafted ‘Initial Prompts’ (line 1 of Alg. 1). These operators undergo “Parallel Operator Evaluation”, receiving scores based on their efficacy across MOPs. These scores inform the “Dynamic Operator Selection” (lines 4-5), bringing diversity in “Crossover Prompts” by altering the number of selected parent operators, thus fostering the creation of innovative operators during “Operator Crossover” (line 6) and “Operator Mutation” (line 7). The newly generated operator obtained through mutation (i.e., $\hat{RC}$) is evaluated in parallel to obtain its score, which is then integrated into the population, followed by an elitist strategy-based update of the population (lines 8-11). The evolution of operators persists, driven by enhanced optimization performance, until the termination criteria are met.
@@ 10       $\mathcal{P}_{ev}=\mathcal{P}_{ev}\cup\{RC\}$
The prompts for initialization are presented in the Fig. 2, titled by “INITIALIZATION”. As depicted, the prompts can be briefly divided into three parts, i.e., ‘System’, ‘Description of Task’ and ‘Requirements’, respectively. The ‘System’ gives the initial prompts for LLM with a brief description for the task. The ‘Description of Task’ outlines the details of the task for the LLM, while the ‘Requirements’ provides the specific requirements on the task. Within these prompts, the placeholder #PROBLEM# is a user-defined descriptor for MOPs; the #PROBLEM_DESC# presents an informative introduction to the MOP with optimization objectives, which aids crafting bespoke search operator for specific MOPs; and the #FORMAT# delineates the operator’s specifics, encompassing the function name, input parameters, output and their annotations. Details of these placeholders have been given in the Supplementary Materials. Employing these comprehensive prompts enables the LLM to generate EA operators that may achieve higher score across MOPs, as delineated in Alg. 2. To further validate the operator produced by the LLM engine, it will undergo Pilot Run & Repair as discussed in Section III-E, specifically lines 3-7 of Alg. 2. The generation process continues until an operator successfully passes the validation tests or a rectified operator is produced through the repair process outlined in Alg. 5.
INITIALIZATION
@@ III-B Dynamic Operator Selection via Scores
Considering the aforementioned issue, this study introduces a dynamic selection mechanism to facilitate the generation of novel operators during operator crossover. Specifically, as illustrated in lines 4-5 of Alg. 1, the selection probabilities for the individual operators within $\mathcal{P}_{ev}$ are determined based on their performance scores on MOPs. A higher score signifies enhanced search capability and, consequently, an increased likelihood of selection. The selection probability for each operator is computed as follows:
@@
where $N_{ev}$ denotes the number of operators within the population. Additionally, a random integer $N_{s}$, where $1<N_{s}\leq N_{max}$, is generated to designate the count of parent operators chosen for the operator crossover.
@@ III-C Operator Crossover
Utilizing the dynamically selected parent operators, this study performs informative operator crossovers by engaging a LLM with these operators. The LLM, serving as an intelligent engine, is anticipated to synthesize innovative operators that surpass the parent operators in terms of scores. The prompts for the LLM are structured as Fig. 3, named by “CROSSOVER”:
CROSSOVER
@@
Figure 3: Prompt design for ‘Operator Crossover’
The placeholders #PROBLEM#, #FORMAT#, #$N_{s}$#, and #SELECTED_OPERATORS# denote the MOP name, the coding format of the operator, the quantity of parent operators selected, and the XML code snippets of the parent operators with their scores, respectively. The dynamic nature of these prompts is likely to facilitate the generation of varied operators by the LLM, as outlined in line 3 of Alg. 3. However, as outlined in line 4 of Alg. 1, the resultant operator $TC$ requires additional validation or refinement to obtain $RC$, akin to the proposed operator initialization.